OpenStudy (anonymous):

In order to solve the following system of equations by addition, which of the following could you do before adding the equations so that one variable will be eliminated when you add them? 4x-2y=7 3x-3y=15 A.Multiply the top equation by 3 and the bottom equation by 4 B.Multiply the top equation by 3 and the bottom equation by 2 C.Multiply the top equation by -3 and the bottom equation by 2 D.Multiply the top equation by 1/3.

OpenStudy (anonymous):

try typing out each of those cases

OpenStudy (anonymous):

for example (A) 12x - 6y = 21 12x - 12y = 60

OpenStudy (anonymous):

so we replace the x and y ?

OpenStudy (anonymous):

no multiply all the numbers by the given value

OpenStudy (anonymous):

and once you've done that, how can you figure it out?

OpenStudy (anonymous):

you add the two values with an "x" next to them together and the two values with a "y" next to them and hope you get 0. If you dont get 0 then that choice is wrong

OpenStudy (anonymous):

so say A) 12x+12 and -6y+-12y and then you hope that both equations equal zero?

OpenStudy (anonymous):

A) 12 + 12 = 24 not 0 -6 -12 = -18 not 0 hence option A is wrong

OpenStudy (anonymous):

oh okay so both equations have to equal to 0

OpenStudy (anonymous):

no just one

OpenStudy (anonymous):

oh okay I get it now